OpenStudy (anonymous):

Solve the equation. 4 over x^2-2x-35 = 5 over x^2+9x+20=

OpenStudy (anonymous):

4 over x^2-2x-35 = 5 over x^2+9x+20= 4 /( x^2-2x-35) = 5/(x^2+9x+20) 5( x^2-2x-35) = 4(x^2+9x+20) 5 x^2-10x-175) = 4x^2+36x+80 combining like terms to get x^2-26x-95=0 now use quadratic fotmula to get x1=29.25 x2=-3.25 approx